OpenStudy (anonymous):

Find the difference negative 1/5 minus negative 3/20. Be sure your answer is reduced to simplest form. PLEASE HELP

OpenStudy (calculusxy):

\[-\frac{ 1 }{ 5 } - (-\frac{ 3 }{ 20 })\]

OpenStudy (calculusxy):

Add -1/5 with the opposite of -3/20. \[-\frac{ 1 }{ 5 } + \frac{ 3 }{ 20 }\]

OpenStudy (calculusxy):

Now you need to follow the rules for adding with different signs. 1. find the absolute value of the addends 2. subtract 3. take the sign of the larger absolute value.